### v1

# Separate Source Channel Coding Is Still What You Need: An LLM-based Rethinking

Abstract
Along with the proliferating research interest in Semantic Communication (SemCom), Joint Source Channel Coding (JSCC) has dominated the attention due to the widely assumed existence in efficiently delivering information semantics. Nevertheless, this paper challenges the conventional JSCC paradigm, and advocates for adoption of Separate Source Channel Coding (SSCC) to enjoy the underlying more degree of freedom for optimization. We demonstrate that SSCC, after leveraging the strengths of Large Language Model (LLM) for source coding and Error Correction Code Transformer (ECCT) complemented for channel decoding, offers superior performance over JSCC. Our proposed framework also effectively highlights the compatibility challenges between SemCom approaches and digital communication systems, particularly concerning the resource costs associated with the transmission of high precision floating point numbers. Through comprehensive evaluations, we establish that empowered by LLM-based compression and ECCT-enhanced error correction, SSCC remains a viable and effective solution for modern communication systems. In other words, separate source and channel coding is still what we need!
Index Terms:
Separate source channel coding (SSCC), joint source channel coding (JSCC), end-to-end communication system, large language model (LLM), lossless text compression, error correction code transformer (ECCT)
I Introduction
Semantic communication (SemCom) has garnered significant attention in recent years, with researchers exploring innovative approaches to enhance the efficiency and reliability of information transmission [1]. Generally, SemCom leverages deep learning-based Joint Source-Channel Coding (JSCC) methods to preserve global semantic information and local texture during the transmission process. DeepJSCC [2] pioneers these works by implementing JSCC with feedback and allowing for real-time adaptation to channel conditions. Along with its steady progress, JSCC has been substantially studied, mostly with the optimization objective shifting from bit error rates to the semantic relevance of the transmitted information in SemCom [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, albeit the awfully exploded research interest, one critical question remains unsolved, that is, why does the joint approach stand out, as a separate source channel coding (SSCC) shall promise more degree of freedom from an optimization perspective?
As the terminology implies, SSCC encompasses two decoupled ingredients, namely source coding and channel coding. The former part lies in effectively compressing the context, and the effectiveness of underlying Deep Neural Networks (DNN)-based predictors, such as Recurrent Neural Networks (RNN)-based DeepZip [15], Long Short Term Memory (LSTM)-based [16, 17] and hybrid DNN-based Dzip [18], toward achieving acceptable text compression, have been validated widely. More prominently, Transformer-based [19] and Large Language Model (LLM)-based compression springs up recently [20, 21, 22, 23, 24]. The latest research [25] unveils the equivalence between compression and prediction. In other words, in the general framework where statistical models predict symbols and encoders use predictive probabilities to perform compression, better predictive models lead directly to better compressors [25]. Hence, the astonishing capability of LLM implies to yield an unprecedented source codec. On the other hand, Error Correction Code (ECC) plays an indispensable role in the channel coding. Although some advanced algebraic block codes like Bose–Chaudhuri–Hocquenghem (BCH) codes [26], Low-Density Parity-Check (LDPC) codes [27] and Polar codes [28], can somewhat ensure the reliability of transmission, the efficient decoding of ECC is an unresolved difficulty. Recently, DNNs start to demonstrate their contribution in channel coding. For example, deep learning models are implemented to achieve Belief Propagation (BP) decoding [29, 30, 31], while a model-free Error Correction Code Transformer (ECCT) for algebraic block codes [32] contributes to the enhancement of the decoding reliability.
In this paper, we introduce an SSCC framework based on LLM and ECCT, and demonstrate the performance superiority over JSCC. This performance improvement primarily arises after tackling the underlying incompatibility between conventional SemCom approaches [3, 7, 8, 9, 10, 11, 12, 13, 14] and digital communication architectures [33]. Particularly, those approaches simply assume the deliverability of encoded semantic feature vectors while neglect the energy cost to transmit high-precision floating point numbers [33]. Unfortunately, further quantization [5, 6] and digital modulation can compromise the widely assumed existence of performance superiority in JSCC. Meanwhile, in contrast to the blunt, direct utilization of the astonishing semantic interpretation capability [34, 35, 36], the deployment of LLMs focuses on the compression and encoding of text to squeeze the largely untapped redundancy. Therefore, our work is also significantly different from those integrating Generative AI (GAI) and SemCom [37, 38, 39, 40, 41, 42, 43]. Furthermore, the adoption of ECCT boosts the effectiveness of SSCC under some cases. In a nutshell, a comprehensive evaluation of LLM and ECCT-based SSCC shows that separate source channel coding is still what we need!
The rest of this paper is organized as follows. Section II introduces the SSCC system model while its key components are enumerated in Section III. Numerical results are presented in Section IV to show the performance superiority of our SSCC system. Finally, Section V concludes this paper with discussions on future works. For convenience, we list the major notations of this paper in Table I.
TABLE I: Major notations used in this paper.
Notation Defination
$\bm{s}_{1:N_{s}},\bm{\hat{s}}$ The transmitted text sequence and the recovered text sequence after source decoding at the receiver side
$\bm{t}_{1:N_{t}}$ The token sequence obtained by tokenizing the text sequence
$C_{s},C_{e}$ The source code and the channel code (error correction code)
$\rho,\tilde{\rho}$ The source distribution and the predicted probability distribution via LLM
$\mathcal{D},D_{i},\tau$ The dictionary of source coder, the $i$-th character in the dictionary and vocabulary of the dictionary
$\mathbb{I}_{k},l_{k},u_{k}$ The probability interval in step $k$ of source coding and its corresponding lower and upper bounds
$\bm{m},\bm{\hat{m}}$ Message encoded by the source coder and the received (and channel decoded) message
$\lambda$ The probability interval, determined by the codeword, in a decimal form
$N,K$ The codeword length and message length of error correction code $C_{e}(N,K)$
$\bm{G,H}$ The generator matrix and the parity check matrix
$\bm{x},\bm{x}_{s},\bm{\hat{x}}$ Codeword encoded by the channel coder, the corresponding sign after BPSK modulation and the soft approximation of transmitted codeword after channel decoding
$\mathcal{N}(\cdot,\cdot),\sigma_{n}$ The Gaussian distribution and the standard deviation of noise
$h$ Channel fading coefficient
$\bm{z},\bm{\tilde{z}},\bm{\hat{z}}$ The additive Gaussian noise, as well as its corresponding multiplicative noise and the prediction result by ECCT
$\bm{y},\bm{y}_{b},\bm{{\tilde{y}}}$ The noisy codeword, its binary form and the result of pre-processing noisy codeword
$\texttt{syn}(\cdot)$ The syndrome of codes defined in ECCT
$f(\cdot)$ The decoding function of ECCT
$\bm{W}$ The learnable embedding matrix for high-dimensional mapping
$\bm{g}(\cdot)$ The code aware self-attention mask
$R_{c},R_{m}$ The code rate and the modulation rate
II System Model
Our SSCC framework encompasses the following ingredients.
  • Source Encoding: The input text sequence denoted as $\bm{s}_{1:N_{s}}$ undergoes a source encoder, which converts characters into a compressed binary message $\bm{m}\in\{0,1\}^{K}$. During source encoding, arithmetic coding (AC) can be leveraged for effective compression here, while for LLM-based processing, an intermediate result (i.e., a sequence of tokens $\bm{t}_{1:N_{t}}$) can be obtained during the transformation from $\bm{s}$ to $\bm{m}$.
  • Channel Encoding and Modulation: The message $\bm{m}$ is then encoded via a LDPC code $C_{e}(N,K)$, which is selected for its excellent error-correction capabilities and compatibility with iterative decoding algorithms, also mentioned in [32]. The encoding process employs a generator matrix $\bm{G}$ to transform the message in $\bm{m}$ to a codeword $\bm{x}\in\{0,1\}^{N}$. The parity check matrix $\bm{H}$, which satisfies $\bm{G}\cdot\bm{H}^{T}=0$ and $\bm{H}\cdot\bm{x}=0$ is a key component of the LDPC decoding process. Afterwards, BPSK modulation maps the binary codeword $\bm{x}$ to a sequence of symbols $\bm{x}_{s}\in\{0,1\}^{N}$ suitable for transmission over the wireless channel. Notably, other error correction codes such as Polar codes [28] can be applied as well.
  • Channel: The modulated signal $\bm{x}_{s}$ is transmitted over a noisy channel, modeled as an Additive White Gaussian Noise (AWGN) channel or a Rayleigh fading channel. The received signal $\bm{y}\in\mathbb{R}^{N}$ is corrupted by additive noise $\bm{z}\sim\mathcal{N}(0,\sigma_{n}^{2})$, resulting in $\bm{y}=h\bm{x_{s}}+\bm{z}$, where $h$ is the channel fading coefficient.
  • Demodulation and Channel Decoding: BPSK demodulation recovers a binary codeword $\bm{x}\in\{0,1\}^{N}$ from $\bm{x}_{s}$. Subsequently, LDPC decoding via the bit-flipping algorithm reconstructs the original message $\bm{m}\in\{0,1\}^{K}$ from $\bm{x}$. Before channel decoding, some complementary procedure such as ECCT can be applied to enhance the decoding performance.
  • Source Decoding: The recovered message $\bm{\hat{m}}$ is ultimately decoded by the source decoder, which reconstructs the original text sequence $\bm{s}$ from the message, effectively reversing the encoding process. Similar to the encoder, the decoder can implement an arithmetic decoding.
In comparison, JSCC often leverages an end-to-end DNN to implement source and channel codecs. Here, the terminology “end-to-end” implies the generally joint training of source and channel codes in most works. The details of JSCC can be found in [1] and the references therein. In the next section, we will address how to leverage the strength of LLM for enhanced text compression and reconstruction, combined with the robustness of ECCT-complemented LDPC codes for error correction, which is shown in Fig. 1.
Figure 1: Framework of LLM-based and ECCT-complemented SSCC system.
III Proposed SSCC Framework
In this section, we introduce LLM-based source coding and ECCT-complemented channel coding.
III-A LLM-based Source Coding
Figure 2: An example of arithmetic coding.
Given the source distribution $\rho$, lossless compression aims to encode a text sequence $\bm{s}_{1:N_{s}}$ sampled from $\rho$ into a binary code $\bm{m}=C_{s}(\bm{s}_{1:N_{s}})$ of minimal possible length $l_{C_{s}}(\bm{s}_{1:N_{s}})$ with no loss of original information. According to Shannon’s source coding theorem [44], the optimal expected bits is $L_{\min}=\mathbb{E}_{\bm{s}\sim\rho}[-\log_{2}\rho(\bm{s})]$. To obtain a code with such optimal length, arithmetic coding [45, 46], a form of entropy encoding, is typically adopted contingent on a probabilistic model over $\rho$ or its marginal distribution. Arithmetic coding implies that frequently used characters be stored with fewer bits while rarely occurred characters correspond to more bits, resulting in fewer bits used in total.
In particular, the input text sequence denoted as $\bm{s}_{1:N_{s}}$ undergoes tokenization by the LLM tokenizer, which converts characters into a sequence of tokens $\bm{t}_{1:N_{t}}$ for LLM to process. The LLM subsequently generates a compact representation of the text, effectively encoding the tokens into a compressed binary message $\bm{m}\in\{0,1\}^{K}$. Specially, considering a dictionary $\mathcal{D}$ of $\tau$ tokens, the input sequence $\bm{s}$ is first parsed into a sequence of $N_{t}$ tokens $t_{1},t_{2},...,t_{N_{t}}$. Given the first $k$ tokens $\bm{t}_{1:k}$, the $(k+1)$-th token $t_{k+1}$ can be inferred as a predicted probability distribution $\tilde{\rho}(t_{k+1}|\bm{t}_{1:k})$. Here, $\tilde{\rho}(t_{k+1}|\bm{t}_{1:k})$ indicates the LLM estimation of the true distribution $\rho(t_{k+1}|\bm{t}_{1:k})$. The incremental decoding nature in LLM implies that it can accurately predict the probability distribution of the next token based on known ones, thus being capable to give a sub-optimal estimation of true distribution [25]. As shown in Fig. 2, selecting the next character is actually narrowing down the probabilistic interval where the sequence locates, which means the code $\bm{m}$ is determined once the interval is fixed. Starting with $\mathbb{I}_{0}=[0,1)$, the previous interval determined by $\bm{t}_{1:k}$ in step $k$ is defined as $\mathbb{I}_{k}=[l_{k},u_{k})$. Therefore, denoting $p(t_{k+1}=D_{j})=\tilde{\rho}(t_{k+1}=D_{j}|\bm{t}_{1:k})$,
$\mathbb{I}_{k+1}(D_{i})=\Big{[}$ $l_{k}+(u_{k}-l_{k})\times\sum\nolimits_{j<i}p(t_{k+1}=D_{j}),$ (1)
$l_{k}+(u_{k}-l_{k})\times\sum\nolimits_{j\leq i}p(t_{k+1}=D_{j})\Big{)}.$
In practice, we consider finite precision arithmetic encoders referring to [47], with description by pseudo-code in Algorithm 1. Consequently, we can find a binary code $\bm{m}=C_{s}(\bm{s}_{1:N_{s}})$ of the shortest length, completely corresponding to the probability interval determined by the sequence. At the receiver side, if the receiver shares a consistent source distribution $\tilde{\rho}$ with sender, given the received (and channel decoded) bit sequence $\hat{\bm{m}}$ corresponding to $C_{s}(\bm{s}_{1:N_{s}})$, we can decode $t_{K_{n+1}}=D_{i}\in\mathcal{D}$ by finding $D_{i}$ such that $\hat{\mathbb{I}}_{n+1}=[l_{n+1},u_{n+1})\subseteq\hat{\mathbb{I}}_{K_{n+1}}(D_{i})=[L,U)$ as in Algorithm 2.
Algorithm 1 Finite-precision arithmetic encoding.
0:  $N_{k}$: Current amounts of emitted bits $\bm{m}_{N_{k}}$
0:  $p_{\text{cum}}(D_{i}|\bm{t}_{1:k})$: Cumulative probability of token $t_{k+1}=D_{i}\in\mathcal{D}$ given the first $k$ tokens
0:  $l_{k},u_{k}$: Current interval determined by the first k tokens
0:  $\varepsilon_{k}$: Number of scaling bits
1:  Initialization:
2:  $N_{k+1}\leftarrow N_{k}$
3:  $l_{k+1}\leftarrow l_{k}+(u_{k}-l_{k})p_{\text{cum}}(D_{i-1}|\bm{t}_{1:k})$ // If $k=0$, use $p_{\text{cum}}(D_{i-1})$
4:  $h_{k+1}\leftarrow l_{k}+(u_{k}-l_{k})p_{\text{cum}}(D_{i}|\bm{t}_{1:k})$ // If $k=0$, use $p_{\text{cum}}(D_{i})$
5:  $\varepsilon_{k+1}\leftarrow\varepsilon_{k}$
6:  Scaling:
7:  while any of the scaling conditions is met do
8:     if $u_{k+1}<0.5$ then
9:         // Scaling 1
10:        $l_{k+1},u_{k+1}\leftarrow 2l_{k+1},2u_{k+1}$
11:        $\bm{m}_{N_{k+1}+1}\leftarrow 0$ // Emit one bit’s 00
12:        $\bm{m}_{N_{k+1}+2:N_{k+1}+1+\varepsilon_{k+1}}\leftarrow 1$ // Emit $\varepsilon_{k+1}$ bits’ $1$
13:        $N_{k+1}\leftarrow N_{k+1}+1+\varepsilon_{k+1}$
14:        $\varepsilon_{k+1}\leftarrow 0$
15:     else if $l_{k+1}\geq 0.5$ then
16:         // Scaling 2
17:        $l_{k+1},u_{k+1}\leftarrow 2(l_{k+1}-0.5),2(u_{k+1}-0.5)$
18:        $\bm{m}_{N_{k+1}+1}\leftarrow 1$ // Emit one bit’s $1$
19:        $\bm{m}_{N_{k+1}+2:N_{k+1}+1+\varepsilon_{k+1}}\leftarrow 0$ // Emit $\varepsilon_{k+1}$ bits’ 00
20:        $N_{k+1}\leftarrow N_{k+1}+1+\varepsilon_{k+1}$
21:        $\varepsilon_{k+1}\leftarrow 0$
22:     else if $0.750.25\leq l_{k+1}<0.5\leq u_{k+1}<$ 0.75 then
23:         // Scaling 3
24:        $l_{k+1},u_{k+1}\leftarrow 2(l_{k+1}-0.25),2(u_{k+1}-0.25)$
25:        $\varepsilon_{k+1}\leftarrow\varepsilon_{k+1}+1$
26:     end if
27:  end while
28:  return  $N_{k+1}$, $\bm{m}_{N_{k}+1:N_{k+1}}$ // Updated emitted bits
Algorithm 2 Finite-precision arithmetic decoding.
0:  $K_{n}$: Current amounts of decoded tokens
0:  $p_{\text{cum}}(D_{i}|\bm{t}_{1:K_{n}})$: Cumulative probability of token $t_{K_{n+1}+1}=D_{i}\in\mathcal{D}$ given the first $K_{n}$ tokens
0:  $l_{n},u_{n}$: Current interval determined by the first $n$ bits
0:  $l_{K_{n}},u_{K_{n}}$: Interval of sequence $\bm{t}_{1:K_{n}}$ that has been decoded
1:  Initialization:
2:  $K_{n+1}\leftarrow K_{n}$
3:  $l_{K_{n+1}},h_{K_{n+1}}\leftarrow l_{K_{n}},h_{K_{n}}$
4:  if the $(n+1)$-th bit $m_{n+1}=0$ then
5:     $l_{n+1},h_{n+1}\leftarrow l_{n},\frac{1}{2}(l_{n}+h_{n})$
6:  else
7:     $l_{n+1},h_{n+1}\leftarrow\frac{1}{2}(l_{n}+h_{n}),h_{n}$
8:  end if
9:  while Not End-of-Sentence symbol do
10:     Search:
11:     Find $D_{i}\in\mathcal{D}$ such that:
12:        $L=l_{K_{n+1}}+(u_{K_{n+1}}-l_{K_{n+1}})p_{\text{cum}}(D_{i-1}|\bm{t}_{1:K_{n}})$ // If $K_{n}=0$, use $p_{\text{cum}}(D_{i-1})$
13:        $U=l_{K_{n+1}}+(u_{K_{n+1}}-l_{K_{n+1}})p_{\text{cum}}(D_{i}|\bm{t}_{1:K_{n}})$ // If $K_{n}=0$, use $p_{\text{cum}}(D_{i})$
14:        $L\leq l_{n+1}<u_{n+1}<U$ // i.e. current interval of the $n$-th bit is included in the interval of $D_{i}$
15:     if $D_{i}$ exists then
16:        Update:
17:        $K_{n+1}\leftarrow K_{n+1}+1$
18:        $t_{K_{n+1}}\leftarrow D_{i}$ // Output $D_{i}$ to the token sequence $\bm{t}$
19:        $l_{K_{n+1}},u_{K_{n+1}}\leftarrow L,U$
20:        Scaling: Similar to the Scaling in Algorithm 1
21:        Go to Search
22:     else
23:        return  $K_{n+1}$, $\bm{t}_{K_{n}+1:K_{n+1}}$
24:     end if
25:  end while
26:  return  $K_{n+1}$, $\bm{t}_{K_{n}+1:K_{n+1}}$ // Updated decoded tokens
Figure 3: LLM-based arithmetic encoding and decoding.
Fig. 3 illustrates such LLM-based arithmetic encoding and decoding where LLM provides a probability interval according to the text sequence $\bm{s}$. Moreover, different from the online setting which trains the model on the data to be compressed, in this paper we assume the availability of the well-trained LLM and employ it to compress different dataset, following the offline setting used in [21].
Remark 1
[25] figure out that the expected code length achieved by leveraging LLM as a compressor could be represented as the cross-entropy, that is,
$H(\rho,\tilde{\rho}):=\mathbb{E}_{s\sim\rho}\left[\sum\nolimits_{i=1}^{n}-\log_{2}\tilde{\rho}(s_{i}|\bm{s}_{<i})\right],$ (2)
where $\rho$ is the source distribution and $\tilde{\rho}$ is the estimation of $\rho$ via a parametric probabilistic model. Hence, the compression shares the same training objective as prediction. Therefore, it can be interpreted as the link between the model log-loss and the compression rate, providing theoretical support for the employment of LLM for source coding.
III-B Error Correction Code Transformer
ECCT [48] belongs to a complementary transformer-alike module to ensure the channel decoding reliability. Notably, ECCT involves specific preprocessing and post-processing steps to effectively avoid overfitting. Without loss of generality, before preprocessing, the syndrome of codes is defined by
$\texttt{syn}(\bm{y}):=$ $\bm{Hy_{b}}=\bm{H}\texttt{sign\_to\_bin}(\bm{y})$ (3)
$=$ $\frac{1}{2}\bm{H}(1-\texttt{sign}(\bm{y}))\in\{0,1\}^{N-K}.$
[…]
$\texttt{sign}(y)=\begin{cases}1,&\quad y>0;\\ 0,&\quad y=0;\\ -1,&\quad y<0.\end{cases}$ (4)
Next, ECCT constructs a $2N-K$ dimensional input embedding by concatenating the element-wise magnitude and syndrome vectors, such that
$\tilde{\bm{y}}:=\big{[}|\bm{y}|,\texttt{syn}(\bm{y})\big{]}\in\mathbb{R}^{2N-K},$ (5)
[…]
The objective of the decoder is to predict the multiplicative noise $\tilde{\bm{z}}$ from $\bm{y}$, where $\bm{y}=h\bm{x}_{s}+\bm{z}=\bm{x}_{s}(h+\bm{x}_{s}\bm{z})=\bm{x}_{s}\bm{\tilde{z}}$. Compared to traditional Transformer architecture [19], ECCT introduces two additional modules for positional reliability encoding and code aware self-attention, as shown in Fig. 4. Note that ECCT handles the channel output $\bm{y}$ as input and provides a prediction $\hat{\bm{z}}$ of the multiplicative noise $\tilde{\bm{z}}$, while its major differences over traditional Transformer architectures are highlighted in the red box of Fig. 4.
Figure 4: The ECCT architecture.
Algorithm 3 Pseudo code of building the attention mask.
0:  parity-check matrix $\bm{H}$ of error correction code $C_{e}(N,K)$
1:  $\text{mask}\leftarrow\texttt{eye}(2N-K)$
2:  for $i=1,2,...,N-K$ do
3:     $\text{idx}\leftarrow\text{where}(\bm{H}[i]==1)$
4:     for $j$ in idx do
5:        $\text{mask}[N+i,\text{ }j],\text{ }\text{mask}[j,\text{ }N+i]\leftarrow 1$
6:        for $l$ in idx do
7:           $\text{mask}[j,\text{ }l],\text{ }\text{mask}[l,\text{ }j]\leftarrow 1$
8:        end for
9:     end for
10:  end for
[…]
12:  return  mask // Output attention mask $\bm{g(H)}$
  • For the channel output $\bm{y}$, the positional reliability encoding transforms each dimension of $\bm{\tilde{y}}$ into a high d dimensional embedding $\bm{\phi}$, which enriches the information of input embedding vectors and replaces $\bm{\tilde{y}}$ as the input of ECCT, defined by
    $\bm{\phi}_{i}=\begin{cases}|\bm{y}_{i}|\bm{W}_{i},&\text{if }i\leq n;\\ \texttt{bin\_to\_sign}\big{(}\texttt{syn}(\bm{y}_{i-n+1})\big{)}\bm{W}_{i},&\text{otherwise}.\end{cases}$ (6)
    where $\{\bm{W}_{i}\in\mathbb{R^{d}}\}_{i=1}^{2n-k}$ denotes the learnable embedding matrix representing the bit’s position dependent one-hot encoding. The encoding method corresponds to the input reliability and is positional, since unreliable information of low magnitude would collapse to the origin, while the syndrome would scale negatively, hence the name positional reliability encoding.
  • The code aware attention mask mechanism aims to integrate code-specific sparse marks which incorporate the inherent structural characteristics of their respective PCM as the domain knowledge. Given a codeword defined by matrix $\bm{G}$ and $\bm{H}$, the attention mask is defined by $\bm{g(H)}:\{0,1\}^{(n-k)\times n}\rightarrow\{-\infty,0\}^{(2n-k)\times(2n-k)}$, the construction of which is shown in Algorithm 3. Then the code aware self-attention mechanism could be represented as
    $A_{H}(\bm{Q,K,V})=\text{Softmax}\left(\frac{\bm{QK}^{T}+\bm{g(H)}}{\sqrt{d}}\right)\bm{V}.$ (7)
    During the implementation the code aware attention mask mechanism is used as an enhancement of the multi-head-attetion module in the classical transformer architecture.
Finally, the training process aims to minimize the Binary Cross Entropy (BCE) loss between predicted noise $\hat{\bm{z}}$ and the multiplicative noise $\tilde{\bm{z}}$, given by
loss $=\texttt{BCELoss}(\hat{\bm{z}},\tilde{\bm{z}})$ (8)
$=-\frac{1}{N}\sum\nolimits_{i}\Big{(}\texttt{bin}(\tilde{\bm{z}}_{i})\cdot\log\big{(}\sigma(\hat{\bm{z}}_{i})\big{)}$
$\quad+\big{(}1-\texttt{bin}(\tilde{\bm{z}}_{i})\big{)}\cdot\log\big{(}1-\sigma(\hat{\bm{z}}_{i})\big{)}\Big{)},$
where $\sigma(\cdot)$ denotes the sigmoid activation function.
Remark 2
The estimation of multiplicative noise is represented as $\hat{\bm{z}}=f(\tilde{\bm{y}})$, while the post-processing step estimates $\bm{x}$ by $\hat{\bm{x}}=\texttt{sign\_to\_bin}(\bm{y}\cdot f(\tilde{\bm{y}}))$. Given that, for correct estimation, $\texttt{sign}(\hat{\bm{z}})=\texttt{sign}(\tilde{\bm{z}})$. Therefore,
[…]
In other words, ECCT contributes to noise-free channel coding.
IV Experiments
In this section, we compare the proposed method with traditional SSCC approaches and existing JSCC solutions under both AWGN and Rayleigh fading channels.
IV-A Simulation Settings
To facilitate comparison, we utilize a pre-processed dataset consisting of the standard proceedings of the European Parliament [49]. A segment of this dataset is selected as an example and fed as the source to a GPT-2 [50] model for source coding. In this numerical experiment, we primarily choose the smallest gpt2-base model with $124$ million parameters while larger models (e.g., $355$-million-parameter gpt2-medium, $774$-million-parameter gpt2-large, and $1.5$-billion-parameter gpt2-xl) are subsequently used for comparative analysis. The arithmetic coding based on the LLM is configured with a precision limit of $31$ bits. For channel coding, we adopt an LDPC code with an information word length of $24$ and a codeword length of $49$, denoted as $\text{LDPC}(49,24)$, resulting in a code rate close to $1/2$. Subsequently, ECCT is used for algebraic block code decoding, which is capable of training on diverse error correction codes. The hyperparameter settings for ECCT training are detailed in Table II. For comparative analysis, we select DeepSC [3], UT [13], UT with quantization [^1] Compared to DeepSC [3] and UT [13], which directly transmit the encodes floats, UT with quantization maps the encoding results to a fixed number ($30$) of bits for transmission. as benchmark JSCC algorithms. Considering the subsequent Signal-to-Noise Ratio (SNR) performance comparison, both algorithms are trained using mixed precision (i.e., float16), which as discussed later has a minimal negative impact on computing the SNR. Key parameters used for training DeepSC and UT are also listed in Table II. Besides, the traditional approach employs Huffman coding for source coding. Furthermore, BLEU (bilingual evaluation understudy, [51]) and semantic similarity measured by BERT [52] are used to measure the performance, which are widely recognized in natural language processing.
TABLE II: Mainly used hyperparameters in the experiments.
Model Hyperparameter Value
ECCT Learning rate $10^{-4}$
Batch size $128$
Number of decoder layers $6$
Dimension of embedding $32$
Number of attention heads $8$
DeepSC Learning rate $10^{-4}$
Batch size $64$
Number of encoder/decoder layers $4$
Dimension of embedding $128$
Dimension of FFN $512$
Number of attention heads $8$
UT Learning rate $10^{-4}$
Batch size $64$
Number of encoder/decoder layers $3$
Dimension of embedding $128$
Dimension of FFN $1,024$
Number of attention heads $8$
Most existing SemCom works evaluate the performance with respect to the $\text{SNR}=10\log_{10}(\frac{E_{\text{tb}}}{N_{0}})$ (dB), where $E_{\text{tb}}$ denotes the energy associated with transmitting a single bit after source/channel coding and digital modulation, while $N_{0}$ corresponds to the noise power spectral density. However, since different coding and modulation schemes across different communication methodologies result in varying numbers of bits transmitted over the physical channel, such a comparative metric of SNR ignores the differences in delivering different numbers of bits. Instead, referring to the total energy consumption $E_{\text{total}}$ by sending $\text{Num}_{\text{unified}}$ bits through the physical channel in LLM-based SSCC system, we use a consistent definition of SNR, in terms of an LLM-based SSCC reference baseline $\text{SNR}_{\text{unified}}$, as a function of the practical employed bits Num. Mathematically,
SNR $=10\log_{10}(\frac{E_{\text{total}}}{N_{0}\cdot\text{Num}})$
$=10\log_{10}(\frac{E_{\text{total}}}{N_{0}\cdot\text{Num}_{\text{unified}}}\times\frac{\text{Num}_{\text{unified}}}{\text{Num}})$ (10)
$=\text{SNR}_{\text{unified}}+10\log_{10}(\frac{\text{Num}_{\text{unified}}}{\text{Num}}).$
where $\text{SNR}_{\text{unified}}$ is used as an independent variable for aligning $E_{\text{total}}$, while for bit-oriented transmission (resp. float-based JSCC), Num denotes the number of bits (resp. float vectors) transmitted through the channel.
On the other hand, as mentioned in Section I and [33], deep learning-based JSCC systems extract the semantic feature of information to embed vectors in latent space, which is incompatible with digital communication system. For JSCC methodologies like UT [13] and DeepSC [3], transmitting a float number certainly consumes far more energy than delivering a binary bit. In this case, if float16 is adopted, we can roughly assume it consumes another $10\times\log_{10}(16)\approx 12.041$ (dB). Hence, for the float-based JSCC methods, the unified evaluation metric is further modified to maintain a consistent energy consumption across different methodologies. In summary,
SNR (11)
=$=\begin{cases}\text{SNR}_{\text{unified}}+10\log_{10}(\frac{\text{Num}_{\text{unified}}}{\text{Num}})+12.041,&\text{float-based};\\ \text{SNR}_{\text{unified}}+10\log_{10}(\frac{\text{Num}_{\text{unified}}}{\text{Num}}),&\text{otherwise}.\end{cases}$
During evaluation, we perform experiments for different schemes in terms of $\text{SNR}_{\text{unified}}$.
IV-B Numerical Results
In this section, we implement the gpt2-base model as compressor and ECCT-complemented $\text{LDPC}(49,24)$ as the error correction code, and compare it with DeepSC [3], UT [13], UT with quantization [53] and the classical SSCC encompassing Huffman coding and ECCT in Fig. 5. The results clearly demonstrate the performance superiority of the proposed SSCC over the other three schemes.
Figure 5: BLEU and Similarity score versus $\textbf{SNR}_{\textbf{unified}}$ for the same number of transmitted symbols, our proposed LLM-based SSCC, with Huffman coding with $\text{LDPC}(49,24)$ in BPSK; DeepSC, UT and UT with quantization trained over AWGN channel.
Similarly, we evaluate the performance under Rayleigh channel in Fig. 6, the results in which show that our system has a clear advantage in terms of the word-level BLEU score. However, in terms of semantic similarity, both the LLM-based and the traditional Huffman-based SSCC systems exhibit some disadvantages at lower SNRs, but still maintain a noticeable advantage at high SNRs.
Figure 6: BLEU and Similarity score versus $\textbf{SNR}_{\textbf{unified}}$ for the same number of transmitted symbols, our proposed LLM-based SSCC, with Huffman coding with $\text{LDPC}(49,24)$ in BPSK; DeepSC, UT and UT with quantization trained over Rayleigh fading channel.
Figure 7: BLEU-4 score versus SNR for the same number of transmitted symbols, our proposed LLM-based SSCC, with Huffman coding with $\text{LDPC}(49,24)$ in BPSK (without ECCT); DeepSC, UT and UT with quantization trained over AWGN(a) and Rayleigh fading channel(b). Ratio of $E_{\text{total}}$ among different systems(c).
In addition to presenting our key experimental results with $\text{SNR}_{\text{unified}}$ as the alignment metric, in Fig. 7, we provide performance comparisons with traditional SNR alignment, as well as the ratio of $E_{\text{total}}$ used by different systems over our LLM-based solution, so as to illustrate the additional energy consumption of JSCC systems in achieving superior performance. Apparently, JSCC systems in SemCom achieve significant gains mainly due to the extra energy consumption.
Figure 8: BLEU-4 score versus $\textbf{SNR}_{\textbf{unified}}$ for the same number of transmitted symbols, with different code rate using $\text{LDPC}(49,24)$ / $\text{LDPC}(49,30)$ / $\text{LDPC}(49,36)$ in BPSK, compared with situations removing ECCT, over AWGN channel and Rayleigh fading channel.
Afterward, we verify the contributing effectiveness of ECCT [32] by comparing the performance of error correction codes with different coding rates under the same code length. Without loss of generality, the evaluation results on the basis of LDPC under AWGN and Rayleigh fading channel are given in Fig. 8. Notably, while the work in [32] does not include Rayleigh channel results, inspired by the subsequent work on Denoising Diffusion Error Correction Codes (DDECC, [54]), we extend ECCT to Rayleigh channels in a similar manner. It can be observed from Fig. 8 that compared to traditional LDPC decoding methods like bit-flipping, ECCT provides consistent performance improvements. Furthermore, for error correction codes of the same length, lower coding rates demonstrate better recovery of noisy signals under the same SNR. More importantly, without ECCT, traditional algorithms struggle to decode noisy signals under Rayleigh channels effectively, and reducing the coding rate slightly improves the performance trivially. However, ECCT trained on the Rayleigh channel achieves as competitive performance as that under AWGN channels.
Considering the scaling law and emergent abilities of LLMs, we evaluate the performance by combining different models from the GPT-2 family with ECCT-complemented LDPC channel coding (i.e., a high-rate $\text{LDPC}(121,110)$ code). Both the end-to-end SSCC performance in Fig. 9 and the compression rate in Fig. 10 indicate a notable performance improvement after adopting a model larger than gpt2. However, the performance difference among gpt2-medium, gpt2-large, and gpt2-xl is marginal. We hypothesize that while increasing model size beyond a certain threshold contributes significantly to system performance, variations within a specific range of model scales yield diminishing returns.
Figure 9: BLEU and Similarity score of models with versus $\textbf{SNR}_{\textbf{unified}}$ different parameter scales (gpt2, gpt2-medium, gpt2-large, gpt2-xl), using $\text{LDPC}(121,110)$ as the error correction code.
Furthermore, inspired by [55], the performance comparison with Zlib and static Huffman coding in Fig. 10 demonstrates that LLM-based arithmetic coding significantly outperforms traditional methods. Moreover, a scaling law is observed in the compression performance, which somewhat corroborates the findings of [55].
Figure 10: Compression rate comparison between traditional methods (zlib, huffman coding) and LLM-AC.
The experimental results presented in Table III further investigate the influence of token block size on the performance. It can be observed that at higher SNR levels, the performance generally improves as the block size increases, indicating that larger block sizes facilitate enhanced semantic preservation, due to their ability to capture more contextual information. However, at lower SNR levels, the performance declines with an increase in block size, suggesting that smaller blocks may be more resilient to avoid cumulative source decoding error in these challenging scenarios.
TABLE III: The influence of token block size on system performance during LLM-based arithmetic source encoding, for $\text{SNR}=\{-6,0,6\}$.
Block size Similarity BLEU (1-gram) BLEU (4-gram)
$-6$ 00 $6$ $-6$ 00 $6$ $-6$ 00 $6$
$16$ $0.7708$ $0.9157$ $0.9993$ $0.1975$ $0.6452$ $0.9877$ $0.0072$ $0.5081$ $0.9830$
$32$ $0.7123$ $0.9359$ $0.9984$ $0.1725$ $0.5842$ $0.9787$ $0.0055$ $0.4666$ $0.9694$
$64$ $0.7001$ $0.8938$ $0.9999$ $0.1160$ $0.5801$ $0.9969$ $0.0018$ $0.4270$ $0.9922$
$128$ $0.7587$ $0.8573$ $0.9999$ $0.1831$ $0.4344$ $0.9999$ $0.0038$ $0.2529$ $0.9999$
V Conclusion and Discussions
In this paper, we have presented a comprehensive analysis and evaluation of SSCC, in comprehensive comparison with JSCC in the context of SemCom. Our proposed SSCC framework, which integrates Large Language Models (LLMs) for source coding and Error Correction Code Transformer (ECCT) for enhanced channel coding, has shown significant performance improvements over JSCC in terms of recovery performance at the word and semantic levels under both AWGN and Rayleigh fading channels, demonstrating the potential effectiveness of SSCC in information transmission. In particular, through extensive experiments, we have validated the ability to leverage the strong compressive capability of LLMs to eliminate redundancy in text and the robustness of ECCT in enhancing decoding reliability under various channel conditions. In a word, separate source channel coding is still what we need!
Nevertheless, despite the validated performance superiority of SSCC, there remain several important issues worthy clarification and investigation.
  • The performance evaluation of text transmission sounds inspiring. The proposed SSCC framework is channel-agnostic, while given the well-known generality issues, the DNN-based JSCC faces a performance decline when the channel changes significantly. But, an extension to image transmission can be more challenging, and several issues like sequential tokenization await for effective solutions. On the other hand, our experimental experience indicates the accuracy of channel coding is of vital importance for the end-to-end performance. Hence, we only consider a rather low, fixed code rate here. However, a systematic tuning of code rate is also worthwhile.
  • This paper only considers the classical JSCC design, while ignoring the latest quantization and digital modulation techniques that emerge in developing JSCC. For example, [5, 6] show that the utilization of a sparsity module to quantize the image embedding can yield significant performance gain. But, [5, 6] have not compared with the astonishing capability of LLM, and thus it remains unclear whether these amendments will help JSCC surpass LLM-based SSCC under a fair comparison. However, given these inspiring results in this paper, there is no doubt that SSCC shall be carefully improved rather than be discarded.
  • The discussions on JSCC are limited to the scenario to recover the semantics as accurately as possible. For SemCom [1], effectiveness-level or pragmatic communications may target at accomplishing different tasks under remotely controlled, noisy environment, rather than simple recovery of accurate semantics. In those cases, the philosophy behind JSCC may promise independent merits.
  • Extensive works have been done to improve the performance of model-free decoders. For example, [56] proposes a systematic and double mask eliminating the difficulty of identifying the optimal one from numerous PCMs from the same code. For performance enhancement on moderate code-length decoding, U-ECCT is proposed in [57] inspired by U-Net; while in [54], DDPM [58] is employed to model the transmission over channels as a diffusion process. Furthermore, a foundation model for channel codes is proposed in [59] for application to unseen codes. Therefore, these latest works are worthy being evaluated in the SSCC framework.